# Efficient Algorithm Level Error Detection for Number-Theoretic Transform Assessed on FPGAs

### v1

@@ V Xilinx/AMD FPGA Implementation results
To confirm the efficiency of our presented approach, we chose to assess its performance by applying it to the NTT multiplication involving a 256-degree polynomial. We conducted a benchmark for implementation on different FPGAs, i.e., Xilinx/AMD Zynq Ultrascale+ and Artix-7. The results clearly demonstrate that our proposed efficient error detection schemes maintain a modest overhead while effectively achieving a high level of fault detection. We utilized the High-Level Synthesis (HLS) Vitis development environment to transform our proposed schemes into Register Transfer Level (RTL) hardware descriptions. The generated IP imported to Vivado to report power, utilization, and latency. Tables II and III demonstrate the results of our implementations and the derivations for area, timing, power, and energy. In the Vivado synthesis tool context, area is determined by combining Slices and DSPs, with a conversion ratio where one DSP is deemed equivalent to 100 Slices. Our proposed error detection designs achieve a maximum operational frequency of around 140 MHz across all the FPGAs. Our proposed error detection scheme imposes a maximum overhead of 9% in additional area and introduces a latency increase of up to 13% in clock cycles, at most considering different designs. Our simulation and implementation code is publicly available in our github [^1] https://github.com/KasraAhmadi/NTT_Error_Detection.
V-A Implementation Optimization
@@
Although HLS has shifted the design entry level of abstraction from RTL to C/C++, practical implementation often requires significant source code rewriting to make it HLD-ready. We have carefully taken this into effort as discussed here. This includes the incorporation of pragmas to attain satisfactory performance. Our intended area and timing efforts implementation was realized by strategically inserting the following pragmas into our program.
@@
Table II: AMD/Xilinx Zynq Ultrascale+, xczu4ev-sfvc784-2-i Implementation Results
@@
Table III: AMD/Xilinx Artix-7, xc7a75ti-ftg256-1L Implementation Results
@@
V-A1 Pre-calculate the decoder module values
@@
V-A2 Utilizing task-level pipelining
@@
V-A3 Pipelining the stage function
@@
VI Conclusion
In this paper, we proposed an efficient algorithm level error detection scheme over polynomial multiplication using NTT. We have aimed for low hardware overhead and low latency, suitable for deeply-embedded systems. The presented scheme effectively safeguards cryptographic algorithms that employ NTT multiplication. By performing simulations, we demonstrated that the proposed error detection scheme achieved extensive coverage of faults. Moreover, we implemented our proposed design on two Xilinx/AMD FPGAs. Regarding overhead and latency, the implementation led to minimal additional expenses in hardware (at most 9% area, 13% latency, 7% power, and 13% energy). With high error coverage and acceptable overhead, the proposed schemes in this works are suitable for resource-constrained and sensitive usage models.

### v2

@@ 5.1. Negative Wrapped Convolution
The results clearly demonstrate that our proposed efficient error detection schemes maintain a modest overhead while effectively achieving a high level of fault detection. We utilized the High-Level Synthesis (HLS) Vitis development environment of AMD/Xilinx to transform our proposed schemes into Register Transfer Level (RTL) hardware descriptions. The generated IP imported to Vivado to report power, utilization, and latency. Tables 3 and 4 demonstrate the results of our implementations and the derivations for area, timing, power, and energy. In the Vivado synthesis tool context, area is determined by combining Slices and DSPs, with a conversion ratio where one DSP is deemed equivalent to 100 Slices. Let us first go over Table 4. This table compares the implementation of two designs (our scheme vs. baseline) on the AMD/Xilinx Zynq Ultrascale+ FPGA (xczu4ev-sfvc784-2-i) in terms of area, power, and performance. Our scheme, which includes error detection, uses slightly more hardware resources than the baseline, with higher values for LUTs (1,469 vs. 1,435), FFs (1,188 vs. 1,122), CLBs (343 vs. 340), and DSPs (33 vs. 30). Both designs have similar power consumption (0.46W vs. 0.45W), but our scheme incurs higher latency (3,703 vs. 3,438 clock cycles), total time (26,439 ns vs. 24,547 ns), and energy consumption (12,161 nJ vs. 11,046 nJ) due to the added complexity of error detection. These are reasonable overheads for providing error detection.
Moreover, as seen in Table 5, it summarizes the implementation results of two designs (our scheme vs. baseline) on the AMD/Xilinx Artix-7 FPGA (xc7a75ti-ftg256-1L), comparing their area, power, and timing performance. Our scheme, which includes error detection, uses more resources than the baseline, with higher values for LUTs (1,569 vs. 1,530), FFs (1,693 vs. 1,597), and DSPs (33 vs. 30). Both designs have similar power consumption, but our scheme consumes slightly more (0.2W vs. 0.19W). However, our scheme has increased latency (3,749 vs. 3,482 clock cycles), total time (26,767 ns vs. 24,861 ns), and energy consumption (5,353 nJ vs. 4,723 nJ), while providing fault detection not found in the baseline design. Our proposed error detection designs achieve a maximum operational frequency of around 140 MHz across all the FPGAs. Our proposed error detection scheme imposes a maximum overhead of 9% in additional area and introduces a latency increase of up to 13% in clock cycles, at most considering different designs. Our simulation and implementation code is publicly available in our github [^1] https://github.com/KasraAhmadi/NTT_Error_Detection.
5.1.1. Implementation Optimization
@@
Although HLS has shifted the design entry level of abstraction from RTL to C/C++, practical implementation often requires significant source code rewriting to make it HLS-ready. We have carefully taken this into effort as discussed here. This includes the incorporation of pragmas to attain satisfactory performance. Our intended area and timing efforts implementation was realized by strategically inserting the following pragmas into our program.
@@
Table 3. AMD/Xilinx Zynq Ultrascale+, xczu4ev-sfvc784-2-i Implementation Results
@@
Table 4. AMD/Xilinx Artix-7, xc7a75ti-ftg256-1L Implementation Results
@@
5.1.2. Pre-calculate the decoder module values
@@
5.1.3. Utilizing task-level pipelining
@@
5.1.4. Pipelining the stage function
@@
5.2. Kyber
We implemented our proposed error detection scheme for the official Kyber NIST submission (Round 3) on both x86-64 architecture and a 12th Gen Intel Core i7-12800H 2.4GHz processor, as well as on more resource-constrained devices, such as the 2.4GHz quad-core ARM Cortex-A72 (Raspberry Pi). We utilized the PAPI library to evaluate the performance of the software implementation and measure the overhead introduced by our error detection scheme. The PAPI library provides a standardized interface for accessing performance counters in CPUs, enabling researchers/developers to measure and analyze the performance of applications efficiently. Table 5 demonstrates the results of our implementation using standard Kyber Round 3 implementation as baseline work on 2 different architectures. The code for our simulation and implementation is accessible on our GitHub. On the ARM Cortex-A72, the baseline work requires 8,859 clock cycles to execute NTT operation in Kyber Round 3, whereas our scheme takes 10,316 clock cycles, an increase of approximately 16.5%. This suggests that the additional error detection or enhancements come with a performance cost in terms of clock cycles.
Similarly, on the Intel Core-i7, the baseline work completes in 5,814 clock cycles, while our scheme takes 7,443 clock cycles, an increase of about 28%. The performance overhead on the Intel Core-i7 is higher compared to the ARM v8, indicating that the performance impact of the error detection or other improvements is more significant on this processor.
Table 5. Software Implementation Results For the NTT in Kyber
ARM v8 Cortex-A72 @ 1.5GHz Intel Core-i7 @ 2.4GHz
Baseline work 8,859 clock cycles 5,814 clock cycles
Our Scheme 10,316 clock cycles 7,443 clock cycles
Our Scheme: The design which includes error detection scheme.
5.2.1. Implementation Optimization
According to equations (16) and (17), the overhead introduced by our error detection scheme depends on the encoding and decoding modules. The encoding module involves one addition and one shift operation. For the decoder, we can store coefficients ($\frac{1}{(\frac{\beta}{\omega^{2b(k)+1}}+\alpha)}$ and $\frac{2\beta}{\omega^{2b(k)+1}}$) in memory to eliminate additional computations. The decoding module comprises two multiplications and one addition.
6. Conclusion
In this paper, we proposed an efficient algorithm level error detection scheme over polynomial multiplication using the NTT and Negative Wrapped Convolution and the NTT operation in Kyber Round 3. We have aimed at achieving low hardware overhead and low latency, suitable for deeply-embedded systems. The presented scheme effectively safeguards cryptographic algorithms that employ the NTT multiplication. By performing simulations, we demonstrated that the proposed error detection scheme achieved extensive coverage of faults. Moreover, we implemented our proposed error detection schemes on two Xilinx/AMD FPGAs and two CPUs. Regarding overhead and latency, the implementation led to minimal additional expenses in hardware and software. With high error coverage and acceptable overhead, the proposed schemes in this work are suitable for resource-constrained and sensitive usage models.